OpenStudy (anonymous):

Gary is using a combination of regular polygonal tiles for his kitchen floor. He begins the tessellation at a vertex by placing one square and two triangles. Which two polygons should he use to fill up the vacant space at the vertex and make it possible to tessellate the floor?

OpenStudy (anonymous):

2 hexagons because 360 – (120 + 60 +60) = 120 2 squares because 180 = 90 + 90 1 triangle and 1 hexagon because 180 = 60 + 120 1 square and 1 triangle because 360 – (90 + 60 + 60) = 90 + 60

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

it would help if i could make out a picture for it

OpenStudy (anonymous):

There's no picture..... :(

OpenStudy (amistre64):

"by placing one square and two triangles" i got no idea what to do with such a vague statement :/

OpenStudy (anonymous):

me neither..... :(

OpenStudy (anonymous):

@Hero can you figure this out?

OpenStudy (amistre64):

http://openstudy.com/study#/updates/4f3c1f72e4b0fc0c1a0f48c4 this question has been batted around before

OpenStudy (amistre64):

if James cant even figure it out ... there is no hope for us mere mortals :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

should I guess?

OpenStudy (amistre64):

thats up to you.

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

http://openstudy.com/study#/updates/4ff7684fe4b01c7be8ca3910 this ones got a semblance of something to it